# Kelvin waves in nonequilibrium universal dynamics of relativistic scalar field theories

### v1

@@ IV.5 Self-similar scaling and decay width
In Ref. [31] it was shown that the transport/kelvon peak in both $F(t,\omega,p)$ and $\rho(t,\omega,p)$ exhibits self-similar scaling in the three-dimensional $\mathrm{O}(1)$ theory. Similarly, in Ref. [122], self-similar scaling was demonstrated for the three-dimensional $\mathrm{U}(1)$ theory for the same peak in $F(t,\omega,p)$ with the exponents $\beta\approx 1/2$ and $z=2$, and for the Bogoliubov peak in $\rho(t,\omega,p)$, with the same $\beta$ and $z\approx 1$.
@@
The decay width of the kelvon peak as extracted from $F$ using the fit form (25) is shown in Fig. 11 for two (left panel) and three (right) dimensions as a function of momentum at different times. In three dimensions, the decay width was found to be time-independent for low momenta and to have a quadratic form $\sim p^{2}$ [31], as we confirm in the right panel. In contrast, as shown in the left panel, it depends on time in a nontrivial way in two dimensions. One finds a relatively large plateau, i.e., an approximately constant decay channel, in addition to a momentum-dependent decay width that decreases with time. This evolution spoils a simple scaling ansatz in $F(t,\omega,p)$.
We finally note that in both $d=2$ and $d=3$ dimensions there is a finite decay width even at zero momentum. However, it is worth noting that in $d=2$, this offset seems to dominate $\gamma_{\mathrm{K}}(p)$ over a vast momentum region, while in $d=3$, the decay width is dominated by a quadratic form $\sim p^{2}$ already at low momenta, facilitating the self-similar scaling solution found in [31, 122].
@@ Acknowledgements.
The authors thank J. Berges, A. N. Mikheev, A. Piñeiro Orioli, D. Proment, I. Siovitz, and A. Villois for discussions on the topics described here. This work is part of and funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1–390900948 (the Heidelberg STRUCTURES Excellence Cluster) and the Collaborative Research Centre, Project-ID No. 273811115, SFB 1225 ISOQUANT. This research was funded in part by the Austrian Science Fund (FWF) [10.55776/P34455]. The authors acknowledge support by the state of Baden-Württemberg through bwHPC (bwUniCluster 2.0).
@@ Appendix A Vortex (line) density detection algorithm
Vortex defects are characterised by phase singularities in $\mathrm{arg}(\phi+\mathrm{i}\pi)$ where a sharp minimum in the energy density also occurs on the lattice. In two dimensions, we compute the vortex density by counting the total number of (anti)vortices in the phase $\mathrm{arg}(\phi+\mathrm{i}\pi)$, and divide this by the (constant) area of the system. First, the input fields $\phi$ and $\pi$ are blockwise averaged over every set of $8\times 8$ lattice points to reduce lattice artefacts. In two dimensions, this process reduces the original $2048^{2}$ lattice to a coarser $256^{2}$ grid. For each point in the phase angle lattice, the algorithm then computes the phase differences around a closed loop. This loop consists of eight neighbouring points (excluding the edges of the lattice) arranged in a clockwise sequence, forming a square path centred around the point of interest. The phase differences between consecutive points in the loop are calculated by subtracting the phase angle at one point from the phase angle at the next point along the loop. Eventually, these differences are summed up and if the total phase change around the loop is approximately $2\pi$, the point is identified as a vortex, and similarly, an antivortex is detected for a phase change of around $-2\pi$. This procedure is designed to avoid mistakenly identifying points as defects where the phase fluctuates between ±$\pm\pi$, as such fluctuations do not result in a uniform phase winding around the point, and therefore do not indicate the presence of a vortex core. Moreover, to avoid duplicate detections, which may still arise due to artefacts, the code enforces a minimum distance between detected vortices and antivortices. After a parameter scan, this minimum distance was chosen as $5$ lattice points. This procedure is then repeated for each point on the $2048^{2}$ lattice.
However, phase singularities in a three-dimensional lattice are more challenging to compute, and doing the computation in the two-dimensional planes, plane by plane, could be misleading. Therefore, we extract the vortex line density [156] from the energy density profiles, e.g. the ones shown in Fig. 2. We first blockwise average over every $6\times 6\times 6$ lattice points to reduce the original $368^{3}$ lattice to a coarser grid. At this point, the energy density becomes quite uniform, and the vortex tangles in Fig. 2 are revealed by taking the points that are at $60\dots 70\%$ of the mean energy value on the lattice. We then compute the number of points that belong to vortex lines, and divide it by the total number of points in the lattice, giving us the density of vortex lines.

### v2

@@ IV.5 Self-similar scaling and decay width
In Ref. Boguslavski and Piñeiro Orioli (2020) it was shown that the transport/kelvon peak in both $F(t,\omega,p)$ and $\rho(t,\omega,p)$ exhibits self-similar scaling in the three-dimensional $\mathrm{O}(1)$ theory. Similarly, in Ref. Piñeiro Orioli and Berges (2019), self-similar scaling was demonstrated for the three-dimensional $\mathrm{U}(1)$ theory for the same peak in $F(t,\omega,p)$ with the exponents $\beta\approx 1/2$ and $z=2$, and for the Bogoliubov peak in $\rho(t,\omega,p)$, with the same $\beta$ and $z\approx 1$.
@@
The decay width of the kelvon peak as extracted from $F$ using the fit form (25) is shown in Fig. 11 for two (left panel) and three (right) dimensions as a function of momentum at different times. In three dimensions, the decay width was found to be time-independent for low momenta and to have a quadratic form $\sim p^{2}$ Boguslavski and Piñeiro Orioli (2020), as we confirm in the right panel. In contrast, as shown in the left panel, it depends on time in a nontrivial way in two dimensions. One finds a relatively large plateau, i.e., an approximately constant decay channel, in addition to a momentum-dependent decay width that decreases with time. This evolution spoils a simple scaling ansatz in $F(t,\omega,p)$.
We finally note that in both $d=2$ and $d=3$ dimensions there is a finite decay width even at zero momentum. However, it is worth noting that in $d=2$, this offset seems to dominate $\gamma_{\mathrm{K}}(p)$ over a vast momentum region, while in $d=3$, the decay width is dominated by a quadratic form $\sim p^{2}$ already at low momenta, facilitating the self-similar scaling solution found in Boguslavski and Piñeiro Orioli (2020); Piñeiro Orioli and Berges (2019).
@@ Acknowledgements.
The authors thank J. Berges, A. N. Mikheev, A. Piñeiro Orioli, D. Proment, I. Siovitz, and A. Villois for discussions on the topics described here. This work is part of and funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1–390900948 (the Heidelberg STRUCTURES Excellence Cluster) and the Collaborative Research Centre, Project-ID No. 273811115, SFB 1225 ISOQUANT. This research was funded in part by the Austrian Science Fund (FWF) [10.55776/P34455]. The authors acknowledge support by the state of Baden-Württemberg through bwHPC (bwUniCluster 2.0), and by SFB 1225 ISOQUANT as well as Heidelberg University concerning the publication fee.
@@ Appendix A Vortex (line) density detection algorithm
Vortex defects are characterised by phase singularities in $\mathrm{arg}(\phi+\mathrm{i}\pi)$ where a sharp minimum in the energy density also occurs on the lattice. In two dimensions, we compute the vortex density by counting the total number of (anti)vortices in the phase $\mathrm{arg}(\phi+\mathrm{i}\pi)$, and divide this by the (constant) area of the system. First, the input fields $\phi$ and $\pi$ are blockwise averaged over every set of $8\times 8$ lattice points to reduce lattice artefacts. In two dimensions, this process reduces the original $2048^{2}$ lattice to a coarser $256^{2}$ grid. For each point in the phase angle lattice, the algorithm then computes the phase differences around a closed loop. This loop consists of eight neighbouring points (excluding the edges of the lattice) arranged in a clockwise sequence, forming a square path centred around the point of interest. The phase differences between consecutive points in the loop are calculated by subtracting the phase angle at one point from the phase angle at the next point along the loop. Eventually, these differences are summed up and if the total phase change around the loop is approximately $2\pi$, the point is identified as a vortex, and similarly, an antivortex is detected for a phase change of around $2\pi$. This procedure is designed to avoid mistakenly identifying points as defects where the phase fluctuates between $\pm\pi$, as such fluctuations do not result in a uniform phase winding around the point, and therefore do not indicate the presence of a vortex core. Moreover, to avoid duplicate detections, which may still arise due to artefacts, the code enforces a minimum distance between detected vortices and antivortices. After a parameter scan, this minimum distance was chosen as $5$ lattice points. This procedure is then repeated for each point on the $2048^{2}$ lattice.
However, phase singularities in a three-dimensional lattice are more challenging to compute, and doing the computation in the two-dimensional planes, plane by plane, could be misleading. Therefore, we extract the vortex line density Vinen (1957) from the energy density profiles, e.g. the ones shown in Fig. 2. We first blockwise average over every $6\times 6\times 6$ lattice points to reduce the original $368^{3}$ lattice to a coarser grid. At this point, the energy density becomes quite uniform, and the vortex tangles in Fig. 2 are revealed by taking the points that are at $60\dots 70\%$ of the mean energy value on the lattice. We then compute the number of points that belong to vortex lines, and divide it by the total number of points in the lattice, giving us the density of vortex lines.